OpenStudy (anonymous):

Which of the following is a factor of x2+ 20x + 96? x + 4 x + 6 x + 3 None of the above

OpenStudy (anonymous):

\[x^{2}+20x+96 = (x+12)(x+8)\] So None of the above!

OpenStudy (e.cociuba):

x 2 + 20x + 96 = x 2 + 12x + 8x + 96 = x(x +12) +8(x+12) = (x+8)(x +12) x +8 and x+12 are factors So the ANSWER: none of the above Credit given to @Soumyadeep

OpenStudy (anonymous):

@e.cociuba much better explanation than mine! :)

OpenStudy (e.cociuba):

Haha u Gave one first, so ya.. lol